# Heterogeneous Trader Responses to Macroeconomic Surprises: A Simulation of Order‑Flow Dynamics

### v1

@@ 3.2 Events Expectations and Surprises
Before scheduled economic announcements, platforms like Bloomberg post publish consensus forecasts. For example, if the market expects the next month’s CPI to be 3.3%. And if the actual CPI is 3.5%, the surprise, defined as the difference between the actual and the expected value, will be
A good surprise means the data is better than expected. However, in the case of CPI, a higher CPI implies higher interest rates, reduced spending, and an increased cost of borrowing; hence, a positive surprise number is actually considered a bad surprise.
@@ 3.3 Asset Classification
In the DGP, we assume to have two types of asset. Risk Free Asset; Risk On Inflation Sensitivity Asset.
@@ 3.4 Extended CAPM Framework
Here, is the traditional market premium while measures the asset’s sensitivity to overall market movements. captures the asset’s response to the overall CPI surprise. And measures the permanent component of the CPI surprise. And notice there might be structural changes in the market condition after some huge shocks, we then decide to use to measure how much of the asset’s return is affected by these long lasting changes, ie. huge inflation rate that could alter the entire market. And the here accounts for asset-specific characteristics not captured by other factors.

### v2

@@ 3.1 Market Participants Heterogeneity
We classify traders into four archetypes—Retail, Pension Funds, Institutional Investors, and Hedge Funds—each with fixed risk aversion levels and distinct information sets. This stratification is consistent with the empirical finance literature, where market participants are observed to differ systematically in both preferences and sophistication:
Retail investors: Typically more risk-averse, with limited access to information and execution technologies. Empirical evidence (e.g., Zhu 2023, JPMorgan 2023) shows retail flows are more dispersed, less sensitive to fundamentals, and often contrarian around announcements.
@@ 3.2 Events Expectations and Surprises
Scheduled macroeconomic announcements, such as CPI releases, are preceded by consensus forecasts published by data providers like Bloomberg and Refinitiv. Markets react not to the level of the data but to the surprise component—defined as the difference between realized and expected values:
For example, if CPI is expected at 3.3% and arrives at 3.5%, the surprise is +0.2 percentage points. Importantly, the sign of a surprise is state-dependent. For instance positive CPI surprise in this case signals stronger inflationary pressure, raising the probability of tighter monetary policy, higher yields, and weaker equities.Conversely, a negative CPI surprise suggests easing inflationary pressure, supporting risk assets but hurting inflation hedges.
@@ 3.3 Asset Classification
For tractability, the DGP begins with two representative assets one is Risk-free asset where offering stable returns and in this case insensitive to interest rate surprises like a short term treasury. And another is Inflation-sensitive risky asset: A stylized equity or bond index that reacts directly to inflation news.This binary setup balances simplicity with interpretability: traders choose between a safe allocation and an asset whose payoff is linked to macroeconomic conditions. While real markets feature a richer asset space (equities, bonds, commodities, derivatives), a two-asset system suffices to highlight cross-sectional differences in trader behavior and utility-driven rebalancing.
@@ 3.4 Extended CAPM Framework
Here, is the traditional market premium while measures the asset’s sensitivity to overall market movements and often note as market premium. captures the asset’s response to the overall CPI surprise. And measures the permanent component of the CPI surprise. In this case it will be the shock premium. And notice there might be structural changes in the market condition after some huge shocks, we then decide to use to measure how much of the asset’s return is affected by these long lasting changes, ie. huge inflation rate that could alter the entire market. And the here accounts for asset-specific characteristics not captured by other factors.